OpenStudy (anonymous):

can someone explain how you turn .1428 into 1/7

OpenStudy (bibby):

\[.1428 = \frac{ 1 }{ x }\]

OpenStudy (anonymous):

So if you get an answer to a problem and the answer is 12.698 how would you find out what fraction that is?

OpenStudy (anonymous):

and putting.698 over 1000 doesn't work

OpenStudy (bibby):

I was gonna say 12 and 698/1000 and then simplify that

OpenStudy (anonymous):

doesnt work. If you put .1428 over 10000 it doesnt give you 1/7. Im so confused

OpenStudy (bibby):

That's because .1428 = 1428/10000

OpenStudy (anonymous):

right..

OpenStudy (bibby):

So what you get when you put .1428 over 10000 is .00001428

OpenStudy (anonymous):

i dont understand how he is getting 1/7

OpenStudy (anonymous):

i understand that one divided by seven is .14 but without know what fraction to use how do you figure it out

OpenStudy (bibby):

Ok, look at it this way. You have a finite decimal and 1/x because you don't know the denominator of the fraction. \[.1428 = \frac{ 1 }{ x }\] \[.1428x = 1\] \[x = \frac{1}{.1428} \]

OpenStudy (bibby):

Do you want me to explain with other examples?